OpenStudy (kainui):

How quickly does the distance between the parabola y=x^2 and y=x^2+1 change as you increase x?

OpenStudy (asnaseer):

by "distance" do you mean the perpendicular distance between the tangent lines at a given value for \(x\)?

OpenStudy (michele_laino):

I think that distace is constant and equal to 1

OpenStudy (michele_laino):

oops...distance...

OpenStudy (asnaseer):

it is constant if you interpret the distance as the vertical distance between the two curves at any given \(x\)

OpenStudy (michele_laino):

I used this formula: \[d=\sup|(f(x)-g(x)|\]

OpenStudy (asnaseer):

or does @Kainui mean the "shortest" distance between the two curves as we increase \(x\)?

OpenStudy (michele_laino):

is the max rate of increasing =1?

OpenStudy (michele_laino):

oops not 1, the formula is: \[\lim _{x \rightarrow +\infty}\frac{ 4 }{ \sqrt{4+(1/x ^{2})} }=\]

OpenStudy (michele_laino):

\[=2\]

OpenStudy (asnaseer):

I think we first need Kainui to clarify what he/she means by "distance" here

OpenStudy (kainui):

By distance I mean shortest length between one parabola to the other.